UT 20131012UT FLAO+PISCES DX

Observer: Arcidiacono,C, Boutsia, K, (INAF),Guerras,JC(LBTO)
Telescope Operator:Bechetti G
SW Support:
AO Support: Guerra,JC
Telescope Support: JUrban
Instrument Support: Edwards,M

Summary

Twice during the night we needed the support of Craig since the commands we sent to PISCES were setting it on a somewhat impossible state and the PISCES was replying with a broken connection, the solution was to restart the IRserver (but I'm not 100% sure since Craig did it for us). Because of that we lost about 1h.
We pointed the core of the globular cluster M15 (we did in 2011) for the calibration of the pixelscale but the preset was failing because the acquisition of the AO star was not converging on the actual star we choose: it was converging on a cluster of stars close-by (~8arcsec). A different and interactive way to acquire the AO star was debugged by Alfio this night and will be used tomorrow.
The AOS stopped to off-load the tilt to the hexapod and this was causing two shell RIP events, we solved this by rebooting the AOS.
Facing the wind increases drastically the force needed to keep the shell position and reduces the forces available range (causing SAFE SKIP and sometimes SHELL RIP) and should be avoided if possible.

Project_RUN Objects Filter ExpTime Notes
McCarthy GJ1215 (K, H) + ND 200s, 150sec COMPLETED - the first 100s in Ks might have a slightly worse PSF
Cusano 652345526 J 420sec REPEAT - We had to abort the script because when we started dither pattern we were too low in elev and the shell ripped
McCarthy BD30_3639 H2, Br-g 330s, 300sec COMPLETED - sky frames also obtained - Br-g images are close to saturation
Cusano FL_Tau Ks, H 358s, 353s This target is COMPLETED - for the program is 20%
DDT XXX H 10min To be repeated on 20131014 - 15min of images but some are with the loop open or poor AO correction - they could be used for sky
Cesaroni IRAS_05137 Ks, H2 730s, 1800 The Ks and H2 filtrers are COMPLETED - for the program 70%

Detail

01:52 We are ready - pointing correction done - collimation done
  • also a test AO preset done on a nearby AO star AO713
  • loop closed with bin4 - OK

01:54 slew to McCarthy - GJ1215
  • guide star not found - check pointing again

10:59 send preset again - seeing 1.2"-1.3" on the GCS
  • AO failed - when focus position changed
  • repeat preset when shell is set
02:08 preset GJ1215 again

02:20 dither pattern for GJ1215 - Ks - 10s exp
  • seeing on GCS 1.1" - DIMM 1.3"

02:25 dither pattern for GJ1215 - H - 15s exp
  • problems with pisces server - we cannot connect

02:36 we contact DonMcCarthy

02:42 we cancel preset - go to zenith - Elliot goes up to see if everything is on for pisces

02:55 psices seems to be working after Craig's intervention
  • we slew to the same target again

03:04 dither pattern in H band - 15s exptime
  • repeat also K because counts during first dither may be a little low
  • dither pattern in K band - 10s exptime

03:18 slew to CUS_652345526 - with ND
  • the shell ripped! - failed during "optimize gain"
  • send preset again
  • seeing GCS: 1.0" - DIMM 1.1"

03:37 the source is too faint for the ND - with 60s in Ks, we get 3000counts
  • cancel preset - telescope to zenith to get the ND out

03:46 we slew to CUS_652345526 again with no ND
  • change focus position in the idl for the NON-ND position
  • repeat runao with correct focus positions

04:00 dither pattern short exposures - J/K
  • during the first offset the AO loop opens - skip frames during the offset..

04:08 need to send the preset again...

04:15 start dither pattern JKKJ
  • we are at 30elev
  • the shell ripped beacause we were too low - we only managed to get the J images
  • repeat this target tomorrow
  • some frames obtained outside script (because pisces had to finish the dither pattern) - these could be used for sky subtraction

04:30 slew to BD30_3639

04:40 start dither pattern - shell ripped!!!

04:45 send the preset again - runao!

04:47 start dither
  • the ao loop is open - send thd preset again

05:02 send preset again - making dither pattern smaller because it is always failing when it offsets....

05:07 start dither pattern - H2 on source

05:15 start dither pattern - Br-g on source
  • the peak is slightly saturate - 2pixels

05:30 start sky for H2, Br-g, J, H, K, 2.14
  • we obtained sky on more filters in order to obtain short exposures also in those
  • the AO loop open and we don't do the other filters

06:07 slew to M15 - for plate scale calibration and north-south orientation
  • sending an acquire preset because the pointing is off
  • send preset again

06:20 it is difficult to close the loop because the field is overcrowded
  • AO failed
  • need to close loop manually because there are too many stars on the guider

06:34 send preset again
  • run runao alone - trying to close the loop - failed

06:44 send preset again
  • there is no way of closing the AO loop

06:47 slew to Cusano V1041
  • there is only one guide star in the catalog that cannot be acquired due to PA constraints for Pisces
  • we do a track preset
  • still trying presets ACE - TRACK
  • the preset keeps failing

07:10 we slew to FL_Tau
  • problems with the secondary - keeps skipping frames

07:20 preset again FL_Tau

08:10 slew for the XXX target
  • collimate on a nearby star

08:17 set NSIGUI - send the ACE preset
  • the ACE preset failed - we should set an aquire preset

08:20 set NSIGUI - send the ACQUIRE preset
  • the TO centered the source on the guider
  • set NSIGUI again - repeat Acquire preset again
  • failed - the source does not appear on the AO ccd
  • keep trying different methods of centering the source
  • ok now the probe is out of the way - the loop is closed and they will manually set the presets from the AO gui

08:40 start obtaining images after manually offsetting
  • the target is not visible increase exp time to 60s - still nothing

08:50 we see the target now!!!
  • pisces.21031012.10410.fits - .10412.fits OFF
  • pisces.21031012.10413.fits - not correct - discard
  • pisces.21031012.10414.fits - .10416.fits - ON target
  • pisces.21031012.10417.fits - .10419.fits ON target but in different offset positions
  • pisces.21031012.10420.fits - .10422.fits ON target but i
  • pisces.21031012.10423.fits - the loop is open!!

09:11 manually close the AO loop
  • pisces.21031012.10424.fits - .10426.fits - the AO is not very good because the wind is introducing a lot of vibration
  • pisces.21031012.10427.fits - .10428.fits better gain this time - offset position
  • pisces.21031012.10429.fits - the loop is open!!
  • pisces.21031012.10430.fits - .10431,fits - the loop is closed but the AO correction is bad!!
  • we give up on this target since we are facing the wind and the AO correction is really poor

09:50 slew to IRAS_05173

10:07 pisces gave error - telnet sais 'connection refused'

10:19 managed to restart the IRserver for pisces
  • pisces is still not working - though it is reachable now
  • JC calls Craig

10:28 send preset again

10:42 skip frame - AO loop is open
  • need to kill and restart AOS - there is no off-loading

10:47 send preset again - AO loop closed
  • centering the target
  • every time we offset the focus position is changing... (???)

10:59 start with the dither pattern Ks/H2/Ks
  • some Ks sky images without correct header
  • dither pattern in the H2

12:51 the AO loop opened - good Ks images until pisces.21031012.10801.fits
  • from pisces.21031012.10802.fits - 10805.fits - are with the loop open

12:53 test the "click'n'go" procedure to use tomorrow on M15

13:06 start Ks flats - no good - we missed the ramp
  • they start from pisces.21031012.10809.fits

13:40 The dome is closed - start darks for the night
  • 60s, 30s, 25s, 20s

Cut and Paste of the IDL command line

Observe commands and IDL console in the following:

CommandLineCutAndPasteOct12

-- KonstantinaBoutsia - 12 Oct 2013
Topic revision: r7 - 13 Oct 2013, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback